Method of Elimination Solving Problem
13. To solve the following linear system by elimination, Brent firsat multiplied each equation by 10. Explain why he did this step. Complete the solution.
0.3x-0.5y=1.2
0.7x-0.2y=-0.1
Please i am really confused thank you very much..

Answer by stanbon(75887)

Method of Elimination Solving Problem
13. To solve the following linear system by elimination, Brent firsat multiplied each equation by 10. Explain why he did this step. Complete the solution.
0.3x-0.5y=1.2
0.7x-0.2y=-0.1
--------------------
Multiply thru by 10 to get rid of the decimals:
-----------------------------------
3x - 5y = 12
7x - 2y =-10
--------------------
Multiply 1st by 7 and 2nd by 3 to get:
----------------------
21x - 35y = 84
21x - 6y =-30
----------------------
Subtract 1st from 2nd to get:
29y = -66
y = -114/29
-----------------
Substitute in 7x -2y = -10 to solve for "x":
7x -2(-114/29) = -10
7x + (228/29) = (-290/29)
7x = -518/29
x = -74/29
=========================
Cheers,
Stan H.